OpenStudy (anonymous):

How do you fin the sec5pi/4?

OpenStudy (jdoe0001):

\(\bf sec(\theta)\implies \cfrac{1}{cos(\theta)}\qquad thus\qquad sec\left(\frac{5\pi}{4}\right)\implies \cfrac{1}{cos\left(\frac{5\pi}{4}\right)}\implies \square ?\)

OpenStudy (jdoe0001):

or.... check your Unit Circle... what's \(\bf cos\left(\frac{5\pi}{4}\right) ?\)